to estimate the size of the bear population on the Keweenaw Peninsula,conservationist captured,tagged and released 50 bears. One year later,a random sample of 100 bears included only 2 tagged bears.What is the conservationist's estimate of the size of the bear population.
Answer by stanbon(75887)

to estimate the size of the bear population on the Keweenaw Peninsula,conservationist captured,tagged and released 50 bears. One year later,a random sample of 100 bears included only 2 tagged bears.What is the conservationist's estimate of the size of the bear population.
------------------------------------------------------
The tagged bears that are released are like a model population.
The proportion is like this:
bunch/caught = bunch/caught
-------------
[Real pop/50] = [100/2]
# pop = 50(100/2) = 50^2 = 2500
===========================================
Cheers,
Stan H.